OpenStudy (anonymous):

An upscale clothing boutique charges $196 for a brand name pair of jeans, which is 40% more than the amount it costs the store to buy one pair of these jeans. As a benefit of working there, employees can purchase the pair of jeans at 30% off the store's cost. How much (in $) would it cost for an employee to purchase a pair of jeans?

OpenStudy (anonymous):

First you would find the true value of the jeans. \[Price(P) = True Price(T) + (0.40*T)\] \[P = T + (0.40*T)\] Using that formula: P = 196 T = x 196 = x + (0.40*x) 196 = x + 0.4x 196 = 1.4x 140 = x So the store buys the jeans for $140 Since employees get 30% off the store's cost. The formula to find this would be as follows: \[Discount Price(D) = True PriceT - (0.30*T)\] \[D = T - (0.30*T)\] Using this new formula: D = x T = 140 x = 140 - (0.30*140) x = 140 - 42 x = 98 So employees pay $98 dollars for the jeans.